# EasyInv: Toward Fast and Better DDIM Inversion

### v1

@@ Introduction
$\mathbf{z}^{*}_{t-1}=\bar{\alpha}_{t-1}\mathbf{z}^{*}_{t-2}+\bar{\beta}_{t-1}\varepsilon^{*}_{t-1},$ (13)
@@
From Eq. (15), it is evident that $\mathbf{z}^{*}_{t}$ is a weighted sum of $\mathbf{z}_{0}$ and a series of noise terms $\varepsilon^{*}_{i}$. The denoising process of Eq. (1) aims to iteratively reduce the impact of these noise terms. In prior research, the crux of inversion is to introduce the appropriate noise $\varepsilon^{*}_{i}$ at each step to identify a suitable $\mathbf{z}^{*}_{t}$. This allows the model to obtain $\mathbf{z}_{0}$ as the final output after the denoising process. However, iteratively updating $\varepsilon^{*}_{i}$ can be time-consuming, and when the model lacks high precision, achieving satisfactory results within a reasonable number of iterations may be challenging.
@@
where $\eta$ is a trade-off parameter, typically set to $\eta\geq 0.7$. The selection of $\bar{t}$ will be discussed in Sec. Quantitative Results. This approach effectively increases the weight $\mathbf{z}_{0}$ in $\mathbf{z}^{*}_{\bar{t}}$, since:
$\begin{split}\mathbf{z}^{*}_{{\bar{t}-1}}-\mathbf{z}^{*}_{\bar{t}}=(\prod_{i=1}^{\bar{t}-1}\bar{\alpha}_{i})(1-\bar{\alpha}_{\bar{t}})\mathbf{z}_{0}\\ +(1-\bar{\alpha}_{\bar{t}})\sum_{i=1}^{\bar{t}-1}(\bar{\beta}_{i}\prod_{j=i+1}^{\bar{t}-1}\bar{\alpha}_{j})\varepsilon^{*}_{i}+(-\bar{\beta}_{t})\varepsilon^{*}_{\bar{t}}\end{split}$ (17)
Given that $10<\bar{\alpha}_{i}<$ 1, for $i=0\rightarrow\bar{t}$, it follows that $(\prod_{i=0}^{\bar{t}-1}\bar{\alpha}_{i})(1-\bar{\alpha}_{\bar{t}})>0$. Consequently, in comparison to $\mathbf{z}^{*}_{\bar{t}}$, $\mathbf{z}^{*}_{\bar{t}-1}$ carries a higher proportion of $\mathbf{z}_{0}$ and is, therefore, less susceptible to the influence of noise. Our approach, therefore, accentuates the significance of the initial latent state $\mathbf{z}_{0}$, which encapsulates the most comprehensive information regarding the original image, within $\mathbf{z}^{*}_{\bar{t}}$.
Table 1: A comparative analysis of quantitative outcomes utilizing the SD-V1-4 model.
@@
Table 2: A comparative analysis of half- and full-precision EasyInv utilizing the SD-V1-4.
@@
Experimentation
We compare our EasyInv over the vanilla DDIM Inversion  (Couairon et al. 2023), ReNoise (Garibi et al. 2024), Pan et al.’s method (Pan et al. 2023) (referred to as Fixed-Point Iteration), using SD V1.4 and SD-XL on one NVIDIA GTX 3090 GPU.
@@
For Fixed-Point Iteration (Pan et al. 2023), we re-implemented it using settings from the paper, as the source code is unavailable. We set the data type of all methods to float16 by default to improve efficiency. The inversion and denoising steps $T=50$, except for Fixed-Point Iteration, which recommends $T=20$. For our EasyInv, we use $0.85\cdot T<\bar{t}<0.95\cdot T$ and $\eta=0.8$ with the SD-XL framework, and $0.05\cdot T<\bar{t}<0.25\cdot T$ and $\eta=0.5$ with SD-V1-4, due to the varying capacities of the two models.
@@
For quantitative comparison, we use three major metrics: LPIPS index (Zhang et al. 2018), SSIM (Wang et al. 2004), and PSNR. The LPIPS index uses a pre-trained VGG16 (Simonyan and Zisserman 2015) to compare image pairs. SSIM and PSNR measure image similarity. We also report inference time. We randomly sample 2,298 images from the COCO 2017 test and validation sets (Lin et al. 2014). With the well-trained SD-XL model, error accumulation is minimal, making all methods perform similarly. Therefore, we display results using the SD-V1-4 model.
Figure 3: A visual assessment of various inversion techniques utilizing the SD-XL model.
Figure 4: A visual assessment of various inversion techniques utilizing the SD-V1-4 model.
Quantitative Results
@@
Table 1 presents the quantitative results of different methods. EasyInv achieves a competitive LPIPS score of 0.321, better than ReNoise (0.316) and Fixed-Point Iteration (0.373), indicating closer perceptual similarity to the original image. For SSIM, EasyInv achieves the highest score of 0.646, showing superior structural similarity crucial for maintaining image coherence. For PSNR, EasyInv scores 30.189, close to ReNoise’s highest score of 31.025, indicating high image fidelity. EasyInv completes the inversion process in the fastest time of 5 seconds, matching DDIM Inversion, and significantly quicker than ReNoise (16 seconds) and Fixed-Point Iteration (14 seconds), highlighting its efficiency without compromising on quality. In summary, EasyInv performs strongly across all metrics, with the highest SSIM score indicating effective preservation of image structure. Its efficient inversion makes it highly suitable for real-world applications where both quality and speed are crucial.
@@
Figure 5: More visual results of our EasyInv utilizing the SD-V1-4 model.
Qualitative Results
We visually evaluate all methods using SD-XL and SD-V1-4. Figure 3 presents a comparison of several examples across all methods utilizing SD-XL. ReNoise struggles with images containing significant white areas, resulting in black images. The other two methods also perform poorly, especially evident in the clock example. Figure 4 displays the results obtained from the SD-V1-4 using images sourced from the internet. These images also feature large areas of white color. ReNoise consistently produces black images with these inputs, indicating an issue inherent to the method rather than the model. Fixed-Point Iteration and DDIM Inversion also fail to generate satisfactory results in such cases, suggesting these images pose challenges for inversion methods. Our method, shown in the figure, effectively addresses these challenges, demonstrating robustness and enhancing performance in handling special scenarios. These findings underscore the efficacy of our approach, particularly in addressing challenging cases that are less common in the COCO dataset.
Figure 5 presents more visual results of our method, with original images exclusively obtained from the COCO dataset (Lin et al. 2014). The results are unequivocal: our approach consistently generates images that closely resemble their originals post-inversion and reconstruction. The variety of categories represented in these images underscores the broad applicability and consistent performance of our method. In aggregate, these findings affirm that our technique is not merely efficient but also remarkably robust, adeptly reconstructing images with a high level of precision and clarity.
@@
Figure 6: Results of MasaCtrl (Cao et al. 2023) with prompt “A football”, using inverted latent generated by different methods as input.
Downstream Image Editing
To showcase the practical utility of our EasyInv, we have employed various inversion techniques within the realm of consistent image synthesis and editing. We have seamlessly integrated these inversion methods into MasaCtrl (Cao et al. 2023), a widely-adopted image editing approach that extracts correlated local content and textures from source images to ensure consistency. For demonstrative purposes, we present an image of a “peach” alongside the prompt “A football.” The impact of inversion quality is depicted in Figure 6. In these instances, we utilize the inverted latents of the “peach” image, as shown in Figure 4, as the input for MasaCtrl (Cao et al. 2023). Our ultimate goal is to generate an image of a football that retains the distinctive features of the “peach” image. As evident from Figure 6, our EasyInv achieves superior texture quality and a shape most closely resembling that of a football. From our perspective, images with extensive white areas constitute a significant category in actual image editing, given that they are a prevalent characteristic in conventional photography. However, such features often prove detrimental to the ReNoise method. Thus, for authentic image editing scenarios, our approach stands out as a preferable alternative, not to mention its commendable efficiency.
Limitations
@@
One potential risk associated with our approach is the phenomenon known as “over-denoising,” which occurs when there is a disproportionate focus on achieving a pristine final-step latent state. This can occasionally result in overly smooth image outputs, as exemplified by the “peach” figure in Figure 4. In the context of most real-world image editing tasks, this is not typically an issue, as these tasks often involve style migration, which inherently alters the details of the original image. However, in specific applications, such as using diffusion models for creating advertisements, this could pose a challenge. Nonetheless, our experimental results highlight that the method’s two key benefits significantly outweigh this minor shortcoming. Firstly, it is capable of delivering satisfactory outcomes even with models that may under-perform relative to other methods, as shown in the above experiments. Secondly, it enhances inversion efficiency by reverting to the original DDIM Inversion baseline (Couairon et al. 2023), thereby eliminating the necessity for iterative optimizations. This strategy not only simplifies the process but also ensures the maintenance of high-quality outputs, marking it as a noteworthy advancement over current methodologies.
@@
In conclusion, our research has made significant strides with the introduction of EasyInv. As we look ahead, our commitment to advancing this technology remains unwavering. Our future research agenda will be focused on the persistent enhancement and optimization of the techniques in this paper. This will be done with the ultimate goal of ensuring that our methodology is not only robust and efficient but also highly adaptable to the diverse and ever-evolving needs of industrial applications.
Conclusion
@@
Our EasyInv presents a significant advancement in the field of DDIM Inversion by addressing the inefficiencies and performance limitations in traditional iterative optimization methods. By emphasizing the importance of the initial latent state and introducing a refined strategy for approximating inversion noise, EasyInv enhances both the accuracy efficiency of the inversion process. Our method strategically reinforces the initial latent state’s influence, mitigating the impact of noise and ensuring a closer reconstruction to the original image. This approach not only matches but often surpasses the performance of existing DDIM Inversion methods, especially in scenarios with limited model precision or computational resources. EasyInv also demonstrates a remarkable improvement in inference efficiency, achieving approximately three times faster processing than standard iterative techniques. Through extensive evaluations, we have shown that EasyInv consistently delivers high-quality results, making it a robust and efficient solution for image inversion tasks. The simplicity and effectiveness of EasyInv underscore its potential for broader applications, promoting greater accessibility and advancement in the field of diffusion models.

### v4

@@ 3.4 EasyInv
$\mathbf{z}^{*}_{t-1}=\bar{\alpha}_{t-1}\mathbf{z}^{*}_{t-2}+\bar{\beta}_{t-1}\varepsilon^{*}_{t-1}.$ (18)
@@
From Equation 20, it is evident that $\mathbf{z}^{*}_{t}$ is a weighted sum of $\mathbf{z}_{0}$ and a series of noise terms $\varepsilon^{*}_{i}$. The denoising process of Equation 1 aims to iteratively reduce the impact of these noise terms. In prior research, the crux of inversion is to introduce the appropriate noise $\varepsilon^{*}_{i}$ at each step to identify a suitable $\mathbf{z}^{*}_{t}$. This allows the model to obtain $\mathbf{z}_{0}$ as the final output after the denoising process. However, iteratively updating $\varepsilon^{*}_{i}$ can be time-consuming, and when the model lacks high precision, achieving satisfactory results within a reasonable number of iterations may be challenging.
@@
where $\eta$ is a trade-off parameter. The selection of $\bar{t}$ and $\eta$ will be discussed in Section 4.1. Considering Equation 16, when applying it to solve the inversion problem, we obtain the following equation:
$\mathbf{z}^{*}_{\bar{t}}=K\mathbf{z}^{*}_{\bar{t}}+(1-K)y_{k},$ (22)
where $H$ is ignored since $y_{k}$ and $\bar{x}_{k}$ show be same kind of data. $\mathbf{z}^{*}_{\bar{t}}$ represents the estimated value $\tilde{x}_{k}$. Next, we replace the measured value $y_{k}$ with $\mathbf{z}^{*}_{\bar{t}-1}$, because while $z_{t}$ cannot be directly measured during the inversion process at time-step $t-1$, the difference $v_{t}=z_{t}-z_{t-1}$ is assumed to be Gaussian noise, in accordance with the basic principles of diffusion models. Thus, $z_{t-1}$ can be treated as the measured value $x_{k}$ in Equation 13 for the inversion problem. The equation then becomes:
$\mathbf{z}^{*}_{\bar{t}}=K\mathbf{z}^{*}_{\bar{t}}+(1-K)\mathbf{z}^{*}_{\bar{t}-1}.$ (23)
Typically, determining the Kalman Gain $K$ is the main goal of the Kalman filter algorithm, which can be computationally expensive. However, the approach described in Equation 21 can be viewed as a simplified version of the Kalman filter, where $K$ is treated as a constant value, determined empirically. With this method, the added noise for inverting $\mathbf{z}^{*}_{{\bar{t}-1}}$ to $\mathbf{z}^{*}_{\bar{t}}$ would be:
$\mathbf{z}^{*}_{\bar{t}}-\mathbf{z}^{*}_{{\bar{t}-1}}=\eta(\mathbf{\bar{z}}^{*}_{\bar{t}}-\mathbf{z}^{*}_{{\bar{t}-1}}).$ (24)
Here, $\mathbf{\bar{z}}^{*}_{\bar{t}}$ represents the original $\mathbf{z}^{*}_{\bar{t}}$ on the right-hand side of Equation 21, generated via DDIM inversion in our case. As indicated in Equation 24, the distribution transition from $\mathbf{z}^{*}_{{\bar{t}-1}}$ to $\mathbf{z}^{*}_{\bar{t}}$ in our method mirrors the distribution pattern of DDIM inversion, maintaining the same variance. Due to $\eta<1$, the average value is reduced. In other words, our method preserves the noise pattern of the original method but applies a rescaling factor.
Algorithm 1 Add EasyInv to an existing inversion method
0:  A inversion algorithm $Inv()$, total inversion steps $T$, latent $z$, chosen steps $\bar{t}$, empirical parameter $\eta$
1:  for $t$ in $T$ do
2:     $z_{t+1}$ = $Inv(z_{t},t)$
3:     if $t$ in $\bar{t}$ then
4:        $z_{t+1}$ = $\eta*z_{t+1}+(1-\eta)*z_{t}$
5:     end if
6:  end for
6:  inverted latent $z_{T}$
Algorithm 1 illustrates the approximate operation of our method in conjunction with an inversion framework.
Table 1: A comparative result of quantitative outcomes.
@@
Table 2: A comparative result of half- and full-precision EasyInv.
@@
4 Experimentation
@@
In Table 1 and Table 2, we compare our EasyInv over other inversion methods, using SD V1.4 on one NVIDIA GTX 3090 GPU. For Fixed-Point Iteration (Pan et al., 2023), we re-implemented it using settings from the paper. We set the data type of all methods to float16 by default to improve efficiency. The inversion and denoising steps are $T=50$, except for Fixed-Point Iteration, which recommends $T=20$. For our EasyInv, we set $0.05\cdot T<\bar{t}<0.25\cdot T$ and $\eta=0.5$. In Table 3 we compare the performance of different down-string tasks when using different inversion methods, the dataset and codes we used in this experiment are from PNPinversion (Ju et al., 2024).
@@
We use three major qualitative metrics: LPIPS index (Zhang et al., 2018), SSIM (Wang et al., 2004), and PSNR with the inference time. We sample 2,298 images from the COCO 2017 test and validation sets (Lin et al., 2014).
Table 3: Performance of downstream methods. Results except “Ours+DirectInv” are from PNPinversion (Ju et al., 2024). DDIM inversion (Couairon et al., 2023) for DDIM, Null-text inversion (Mokady et al., 2023) for NT, Negative-Prompt Inversion (Miyake et al., 2023) for NP, StyleDiffusion (Wang et al., 2023) for StyleD, PNPinversion (Ju et al., 2024) for DirectInv, Prompt-to-Prompt (Hertz et al., 2022) for P2P, MasaCtrl (Cao et al., 2023) for MasaCtrl, Pix2Pix-Zero (Parmar et al., 2023) for P2P-Zero and Plug-and-Play (Tumanyan et al., 2023) for PnP.
$\dagger$ averaged results on A800 and RTX3090 since different environment leads to slightly different performance.
* uses Stable Diffusion v1.5 as base model (others all use Stable Diffusion v1.4).
Method Structure Background Preservation CLIP Similarity
Inverse Editing Distance×${}_{\times 10^{3}}$downarrow PSNR $\uparrow$ LPIPS×${}_{\times 10^{3}}$downarrow MSE×${}_{\times 10^{4}}$downarrow SSIM×${}_{\times 10^{2}}$uparrow Whole $\uparrow$ Edited $\uparrow$
DDIM P2P 69.43 17.87 208.80 219.88 71.14 25.01 22.44
NT P2P 13.44 27.03 60.67 35.86 84.11 24.75 21.86
NP P2P 16.17 26.21 69.01 39.73 83.40 24.61 21.87
StyleD P2P 11.65 26.05 66.10 38.63 83.42 24.78 21.72
DirectInv P2P 11.65 27.22 54.55 32.86 84.76 25.02 22.10
Ours+DirectInv P2P 11.58 27.30 53.52 32.37 84.80 24.97 22.00
DDIM MasaCtrl 28.38 22.17 106.62 86.97 79.67 23.96 21.16
DirectInv MasaCtrl 24.70 22.64 87.94 81.09 81.33 24.38 21.35
Ours+DirectInv MasaCtrl 23.82 22.72 87.65 79.73 81.82 24.36 21.32
DDIM P2P-Zero 61.68 20.44 172.22 144.12 74.67 22.80 20.54
DirectInv P2P-Zero 49.22 21.53 138.98 127.32 77.05 23.31 21.05
Ours+DirectInv P2P-Zero 48.02 21.54 136.78 124.00 77.71 23.40 20.93
DDIM PnP 28.22 22.28 113.46 83.64 79.05 25.41 22.55
DirectInv PnP 24.29 22.46 106.06 80.45 79.68 25.41 22.62
Ours+DirectInv PnP 22.88 22.56 102.34 78.57 80.27 25.38 22.53
4.1 Quantitative Results
@@
Table 1 presents the quantitative results of different methods. EasyInv achieves a competitive LPIPS score of 0.321, better than ReNoise (0.316) and Fixed-Point Iteration (0.373), indicating closer perceptual similarity to the original image. For SSIM, EasyInv achieves the highest score of 0.646, showing superior structural similarity crucial for maintaining image coherence. For PSNR, EasyInv scores 30.189, close to ReNoise’s highest score of 31.025, indicating high image fidelity. EasyInv completes the inversion process in the fastest time of 5 seconds, matching DDIM Inversion (Couairon et al., 2023), and quicker than ReNoise (16 seconds) and Fixed-Point Iteration (14 seconds), highlighting its efficiency without compromising on quality. EasyInv performs strongly across all metrics, with the highest SSIM score indicating effective preservation of image structure. Its efficient inversion makes it suitable for real-world applications where both quality and speed are crucial.
@@
In Table 3, we compare the performance of several down-stream tasks using our method and other inversion techniques. Most of the results in this table are from PNPInversion (Ju et al., 2024), where a dataset for 9 different image editing tasks is introduced, along with the corresponding code for both editing and evaluation. We used these codes and dataset for this experiment. The only modification we made was to incorporate our method into DirectInv, the inversion method proposed in their work (Ju et al., 2024), and its results are indicate as ‘ours+DirectInv’ in Table 3. As we pointed out, our method is able to combined with most existing inversion algorithm. The results in Table 3 show the our advantage. By adding our method, the performance of DirectInv improves in 5 out of 7 metrics across all editing tasks, with minimal changes in the remaining 2 metrics.
4.2 Qualitative Results
We visually evaluate all methods using SD-V1-4. Figure 3 displays the results using images sourced from the internet. These images also feature large areas of white color. ReNoise struggles with images containing significant white areas, resulting in black images. Fixed-Point Iteration and DDIM Inversion also fail to generate satisfactory results in such cases, suggesting these images pose challenges for inversion methods. Our method, shown in the figure, effectively addresses these challenges, demonstrating robustness and enhancing performance in handling special scenarios. These findings underscore the efficacy of our approach, particularly in addressing challenging cases that are less common in the COCO dataset. For better illustrations, we show more visual examples in Figure 4.
Figure 3: A visual assessment of various inversion techniques.
@@ 4.2 Qualitative Results
Figure 4: More visual results of our EasyInv.
4.3 Ablation Studies
To clarify our experimental choices, in Table 4 we present ablation studies on the parameter $\eta$, conducted on the same dataset as in Table 3. Note that $\eta=1$ would correspond to a standard DDIM inversion and $\eta=0$ would bypass the inversion operation. Therefore, we exclude them. Our findings indicate that $\eta=0.5$ yields the best overall performance.
Table 4: Ablation Experiments on $\eta$.
Setting Structure Background Preservation CLIP Similarity
$\eta$ Editing Distance×${}_{\times 10^{3}}$downarrow PSNR $\uparrow$ LPIPS×${}_{\times 10^{3}}$downarrow MSE×${}_{\times 10^{4}}$downarrow SSIM×${}_{\times 10^{2}}$uparrow Whole $\uparrow$ Edited $\uparrow$
0.2 PnP 25.90 23.29 113.86 66.11 80.13 24.66 21.60
0.4 PnP 24.44 23.29 107.18 66.32 80.69 24.82 21.76
0.5 PnP 22.88 22.56 102.34 78.57 80.27 25.38 22.53
0.6 PnP 23.36 23.21 101.37 67.96 81.07 25.02 22.00
0.8 PnP 22.89 22.92 98.44 72.30 80.94 25.30 22.29
@@ 4.3 Ablation Studies
Figure 5: Results of MasaCtrl (Cao et al., 2023) with prompt “A football”, using inverted latent by different methods as input.
Figure 6: Comparison from DiffusionTrend (Zhan et al., 2024).
4.4 Downstream Image Editing
To showcase the practical utility of our EasyInv, we have employed various inversion techniques within the realm of consistent image synthesis and editing. We have seamlessly integrated these inversion methods into MasaCtrl (Cao et al., 2023), a widely-adopted image editing approach that extracts correlated local content and textures from source images to ensure consistency. For demonstrative purposes, we present an image of a “peach” alongside the prompt “A football.” The impact of inversion quality is depicted in Figure 5. It is clear that our inversion method demonstrates superior performance in both texture and shape. We also apply our EasyInv to two additional downstream tasks: DiffusionTrend (Zhan et al., 2024) for virtual try-on and UniVST (Song et al., 2024) for video style transferring. In Figure 6, using our method results in better consistency of the cloth compared to DDIM and results in Figure 7 show that DDIM inversion leads to chaotic outcomes, whereas our method yields much clearer results.
Figure 7: Comparison from UniVST (Song et al., 2024).
4.5 Limitations
@@
One potential risk is the phenomenon known as “over-denoising,” which occurs when there is a disproportionate focus on achieving a pristine final-step latent state. This may result in overly smooth image outputs, as exemplified by the “peach” figure in Figure 3. In most real-world image editing tasks, this is not a typical issue, as these tasks often involve style migration, which inherently alters the details of the original image. However, in specific applications, such as using diffusion models for creating advertisements, this could pose a challenge. Nonetheless, our experimental results highlight that the method’s two key benefits outweigh this minor shortcoming. Firstly, it is capable of delivering satisfactory outcomes even with models that may under-perform relative to other methods. Secondly, it enhances inversion efficiency by reverting to the original DDIM Inversion baseline (Couairon et al., 2023), thereby eliminating the necessity for iterative optimizations. This strategy not only simplifies the process but also ensures the maintenance of high-quality outputs, marking it as a noteworthy advancement over current methodologies.
@@
In conclusion, our research has made significant strides with the introduction of EasyInv. As we look ahead, our commitment to advancing this technology remains unwavering. Our future research agenda will be focused on the persistent enhancement and optimization of the techniques. This will be done with the ultimate goal of ensuring that our method is not only robust and efficient but also highly adaptable to the diverse and ever-evolving needs of industrial applications.
5 Conclusion
@@
Our EasyInv addresses the inefficiencies and performance limitations in traditional iterative optimization methods. By emphasizing the importance of the initial latent state and introducing a refined strategy for approximating inversion noise, EasyInv enhances both the accuracy efficiency of the inversion process. Our method reinforces the initial latent state’s influence, mitigating the impact of noise and ensuring a closer reconstruction to the original image. This approach not only matches but often surpasses the performance of existing DDIM Inversion methods, especially in scenarios with limited model precision or computational resources. EasyInv also shows a remarkable improvement in inference efficiency, achieving approximately three times faster processing than standard iterative techniques. Through extensive evaluations, we have shown that EasyInv delivers high-quality results, making it a robust and efficient solution for image inversion tasks. The simplicity and effectiveness of EasyInv underscore its potential for broader applications.
Acknowledgments
This work was supported by the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. U23A20383, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No. 2021J06003, No.2022J06001).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.